# A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics

### v2

@@ 2 Defect chemistry of oxide ceramics
The basics of defect chemistry and the sharp interface bicrystal models (the MS and the GC models) for the equilibrium state were well described in Refs. [30, 40, 16]. Thereby the GB is treated as sharp interface with GB core properties. Thus the concentration profiles of oxygen vacancy and dopant are discontinuous, with concentration jumps across the GB core. For the bicrystal, it becomes a one-dimensional case.
Assume a coordinate system $x$, with the origin $x=0$ located at the GB core. Moreover, point defects are assumed to be dilute and non-interacting in both GB core and bulk phases. Therefore, the electrochemical potentials of oxygen vacancy and dopant in the bulk and in the core region can be formulated [30] as
@@
where $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$, $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$, $\mu_{\text{dop},\text{b}}^{0}$ and $\mu_{\text{dop},\text{c}}^{0}$ denote the standard formation energies of oxygen vacancy and dopant in the bulk and in the core, respectively. The symbols $N_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$, $N_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$, $N_{\text{dop},\text{b}}^{0}$ and $N_{\text{dop},\text{c}}^{0}$ indicate the number of available sites per unit volume of oxygen vacancy and dopant in the bulk and core, respectively. $z_{\text{V}_{\text{O}}^{..}}$ and $z_{\text{dop}}$ denote the charge numbers of oxygen vacancy and acceptor dopant. $\phi(x)$ is the electrostatic potential, and $\Phi_{0}$ is the electrostatic potential in the GB core. Similarly, $c_{\text{dop},\text{c}}$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}$ are the concentrations of dopant and oxygen vacancy in the GB core. Thus, we have $\Phi_{0}=\phi(0)$, $c_{\text{dop},\text{c}}=c_{\text{dop},\text{b}}(0)$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}=c_{\text{V}_{\text{O}}^{..},\text{b}}(0)$. Additionally, $k_{\text{B}}$, $e$ and $T$ are the Boltzmann constant, elementary charge and temperature, respectively. The formation of a SCL requires the segregation of oxygen vacancy in the core. The smaller value of $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$ than $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$, i.e. $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}-\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}<0$, drives the formation of SCL. At the equilibrium state, when electrochemical potentials should be everywhere the same (in other words, no gradient), we have
@@
$\Delta g_{\text{V}_{\text{O}}^{..}}$ The difference of formation energy of oxygen vacancy between core and bulk, i.e. $\Delta g_{\text{V}_{\text{O}}^{..}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}-\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$ eV
$\Delta g_{\text{dop}}$ The difference of formation energy of dopant between core and bulk, i.e. $\Delta g_{\text{dop}}=\mu_{\text{dop},\text{c}}^{0}-\mu_{\text{dop},\text{b}}^{0}$ eV
$N_{\text{V}_{\text{O}}^{..},\text{b}}$ Maximum number density of oxygen vacancy in the bulk phase $\frac{1}{\text{m}^{3}}$
$N_{\text{V}_{\text{O}}^{..},\text{c}}$ Maximum number density of oxygen vacancy in the core region $\frac{1}{\text{m}^{3}}$
$N_{\text{dop},\text{b}}$ Maximum number density of dopant in the bulk phase. $N_{\text{dop},\text{b}}=1/a^{3}$, with $a$ being the perovskite lattice constant $\frac{1}{\text{m}^{3}}$
$N_{\text{dop},\text{c}}$ Maximum number density of dopant in the core region $\frac{1}{\text{m}^{3}}$
$z_{\text{V}_{\text{O}}^{..}}$ Valency number of oxygen vacancy, $z_{\text{V}_{\text{O}}^{..}}=2$
@@ 2.1 The Mott-Schottky model
The Mott-Schottky model assumes that the dopant is immobile and its concentration remains fixed and constant everywhere, i.e. $\nabla c_{\text{dop}}=0$. Note that this assumption simplifies the scenario but breaks the equilibrium condition of the electrochemical potential of the dopants. Based on the assumption, Poisson’s equation can be written as
@@
[…]
@@
[…]
@@ 2.2 The Gouy-Chapman model
The numerical solutions of the MS and GC models have been achieved through an iteration process [30]. In the present work, we propose an alternative numerical method via finite element method (FEM). By considering the symmetric configuration of a bicrysal (ignoring the grain misorientation), the semi-infinite simulation domain of $[0,\infty)$ is sufficient. The location of the GB core is located at the left boundary of the semi-infinite domain ($x=0$). As for the boundary conditions, the electrostatic field is grounded at the right side, i.e., $\phi|_{x=\infty}=0$, while at the left side, a mixed boundary condition is given by
@@
where $Q_{\text{c}}(\phi_{0})=ew_{\text{c}}(z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{c}}+z_{\text{dop}}c_{\text{dop},\text{c}})$ and $\phi_{0}=\phi|_{x=0}$. We employed the boundary element in the finite element method to regard directly the mixed type of boundary condition defined in the above equation. The obtained numerical solutions agree fully with these from Ref. [30] (See Supplementary).
@@ 3.1 Free energy density functional and electrochemical potentials
We include here the defect chemistry into the grain growth model for oxide ceramics, with Fe-doped STO as a prototype. Upon grain contact in Fe-doped STO, the lower formation energy of oxygen vacancy in the core compels point defects to segregate towards the GB core. Afterwards, the depletion or accumulation zone appears in the vicinity of GB core. The SCL forms through a diffusion process. The handling of SCL is similarly addressed in Ref. [30]. The segregation of point defects in the core serves to dissipate the total free energy of the system and also induces solute drag effect when the core moves.
Assuming an Fe-doped STO polycrystalline system consisting of $n$ grains, we choose Kim-Kim-Suzuki (KKS) phase-field model and utilize a set of non-conserved order parameters (OPs) ${\eta_{i}}$ to distinguish between different grains [41]. Within grain $i$, $\eta_{i}=1$ and $\eta_{j}=0$ for $j<n$ and $j\neq i$. Within the core region of grain $i$ and $j$, the order parameters $\eta_{i}$ and $\eta_{j}$ spatially vary from 0 to 1. In the latter context, $\eta$ denotes the entire set of OPs $\eta_{i}$}. When considering electrochemical contributions, we employ two conserved concentration fields along with an electrostatic potential field. Other normalized field variables are $C_{\text{V}_{\text{O}}^{..}}$ for oxygen vacancy concentration, $C_{\text{dop}}$ for dopant concentration, and $\phi$ for the electrostatic potential. In Fe-doped STO, $\text{Fe}^{3+}$ and $\text{Fe}^{4+}$ can coexist. At high temperatures and low oxygen partial pressures, a higher proportion of $\text{Fe}^{3+}$ is observed and thus the dopant $\text{Fe}_{\text{Ti}}^{{}^{\prime}}$ has negative charge. For example, during the sintering process at 1623 K in air, approximately 90% of the iron is estimated to be in the $\text{Fe}^{3+}$ state, according to the bulk defect chemical model (see Fig. A1 in Ref. [42]). Thus, in this study, we assume the valence state of Fe to be 3+ and therefore the valance of dopant $\text{Fe}_{\text{Ti}}^{{}^{\prime}}$ is -1, i.e. $z_{\text{dop}}=-1$. Then, the formulation of total free energy density functional is given by
@@
$f^{\text{grad}}$ represents the gradient term, it is given by
@@
where $\kappa$ is the coefficient of the gradient term. $f^{\text{loc}}$ is the well potential. According to Ref. [43], its formulation is given by
@@
with $\gamma=1.5$ for producing symmetric profile of $\eta$ and isotropic grain growth.
The determination of parameters such as $\kappa$ and $\omega$ can be linked to characteristics such as core width ($w_{\text{c}}$) and interface energy ($\sigma$). Their relations can be expressed as
@@
In the present work, we only consider the isotropic interface energy, i.e. it is not dependent on GB misorientation. The third term in free energy functional is the electrochemical contribution. Here, we treat the electrochemical free energy density as a mixture of bulk phase and core by an interpolation function. Thus, $f^{\text{ech}}$ is
@@
where the subscripts b and c indicate the bulk phase and core, respectively. $h(\eta)$ is the interpolation function and is given by
@@
Here, $h(\eta)$ exhibits non-monotonic behavior. Specifically, for an $i/j$ GB, $h(\eta_{i})=0$ when $\eta_{i}=1$ (representing the bulk phase), whereas $h(\eta_{i})=1$ when $\eta_{i}=0.5$ (representing the core). This interpolation function allows us to differentiate between the bulk phase and the core region. $\epsilon_{0}$ is the vacuum permittivity; $\epsilon_{\text{r}}$ is the relative permittivity. For STO, $\epsilon_{\text{r}}=9\times 10^{4}/(T-35)$ [30]. $\epsilon_{\text{r}}$ is only dependent on temperature in the present work. Hence, $\epsilon_{\text{r}}$ is constant across the SCL zone for specific temperature. Furthermore, $f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})$ and $f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})$ in Eq. 22 denote the electrochemical free energy contributions from the bulk and core, respectively, with $C_{\text{V}_{\text{O}}^{..},\text{b}}$ and $C_{\text{dop},\text{b}}$ being the phase concentrations of oxygen vacancy and dopant in the bulk phase, while $C_{\text{V}_{\text{O}}^{..},\text{c}}$ and $C_{\text{dop},\text{c}}$ being the phase concentrations of oxygen vacancy and dopant in the core. When $h(\eta)=1$, $C_{\text{dop},\text{c}}$ and $C_{\text{V}_{\text{O}}^{..},\text{c}}$ correspond to $c_{\text{dop},\text{c}}$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}$. When $h(\eta)=0$, $C_{\text{dop},\text{b}}$ and $C_{\text{V}_{\text{O}}^{..},\text{b}}$ are same as $c_{\text{dop},\text{b}}(x)$ and $c_{\text{V}_{\text{O}}^{..},\text{b}}(x)$. Local concentration is treated as the mixture of phase concentrations in bulk and core. The relations between local concentration and phase concentration are expressed as
@@
In principle of the thermodynamic model extended by Ref. [30], the expressions for the electrochemical free energies $f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})$ and $f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})$ are given by
@@
$\begin{split}f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}C_{\text{V}_{\text{O}}^{..},\text{c}}+g^{0}_{\text{dop},\text{c}}C_{\text{dop},\text{c}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{c}}\ln(C_{\text{V}_{\text{O}}^{..},\text{c}})+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}})\ln(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}})\right.\\ &\left.\left.+C_{\text{dop},\text{c}}\ln(C_{\text{dop},\text{c}})+(\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}})\ln(\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}})\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{c}}+z_{\text{dop}}C_{\text{dop},\text{c}}\right)\phi,\end{split}$ (26)

### v3

@@ 2 Defect chemistry of oxide ceramics
In this section, we review the classical defect chemistry theory in determining the SCL formation under equilibrium state. The performed defect chemistry calculations are used to benchmark the phase-field simulations. The basics of defect chemistry and the sharp interface bicrystal models (the MS and the GC models) for the equilibrium state were well described in Refs. de2009formation , tong2020analyzing , maier2023physical . According to abrupt GB||SCL model of Jamnik et al. jamnik1995interfaces , the interface region is divided into a GB core and SCL region, in which the materials constants behave as step functions. The two phases (GB core and bulk) are essentially treated as two phases, each with their own properties, and separated by a sharp interface. Thus the concentration profiles of oxygen vacancy and dopant , for $\mu^{0}_{\text{V}_{\text{O}}^{..}}\neq 0$ and $\mu^{0}_{\text{dop}}\neq 0$, are discontinuous, with concentration jumps across the GB core. For the bicrystal, it becomes a one-dimensional case.
Assume a symmetric bicrystal in a coordinate system $x$, with the origin $x=0$ located at the GB core. We introduce $c_{\text{def}}$ point defects into $N_{\text{def}}$ (the subscript def denotes different types of point defects) available number sites in a unit volume. We should note that $N_{\text{def}}$ varies for different sublattices and different phases. When $N_{\text{def}}\gg c_{\text{def}}$, point defects are assumed to be dilute and non-interacting in both GB core and bulk phases. In Ref. maier2023physical , the electrochemical potential of defect is formulated as
$\mu_{\text{def}}=\mu^{0}_{\text{def}}+k_{\text{B}}T\ln\frac{c_{\text{def}}}{N_{\text{def}}-c_{\text{def}}}+z_{\text{def}}e\phi,$ (1)
where $\mu^{0}_{\text{def}}$ is the standard formation energy of defect. $k_{\text{B}}$, $T$, $z_{\text{def}}$, $e$ and $\phi$ are the Boltzmann constant, temperature, valance state of the defect, the elementary charge and the electrostatic potential, respectively. Then, based on the abrupt GB||SCL model aforementioned, the electrochemical potentials of oxygen vacancy and dopant in the bulk and in the core region can be formulated de2009formation as
@@
where $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$, $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$, $\mu_{\text{dop},\text{b}}^{0}$ and $\mu_{\text{dop},\text{c}}^{0}$ denote the standard formation energies of oxygen vacancy and dopant in the bulk and in the core, respectively. The symbols $N_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$, $N_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$, $N_{\text{dop},\text{b}}^{0}$ and $N_{\text{dop},\text{c}}^{0}$ indicate the number of available sites per unit volume of oxygen vacancy and dopant in the bulk and core, respectively. $z_{\text{V}_{\text{O}}^{..}}$ and $z_{\text{dop}}$ denote the charge numbers of oxygen vacancy and acceptor dopant. $\phi(x)$ is the spatially various electrostatic potential in the bulk phase, and $\Phi_{0}$ is the electrostatic potential in the GB core, which taken to be constant. Similarly, $c_{\text{dop},\text{c}}$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}$ are the concentrations of dopant and oxygen vacancy in the GB core. Thus, we have $\Phi_{0}=\phi(0)$, $c_{\text{dop},\text{c}}=c_{\text{dop},\text{b}}(0)$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}=c_{\text{V}_{\text{O}}^{..},\text{b}}(0)$. The formation of a SCL requires the segregation of oxygen vacancy in the core. The smaller value of $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$ than $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$, i.e. $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}-\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}<0$, drives the formation of SCL. At the equilibrium state, when electrochemical potentials should be everywhere the same (in other words, no gradient), we have
@@
$\Delta g_{\text{V}_{\text{O}}^{..}}$
The difference of formation energy of oxygen vacancy between core and bulk,
i.e. $\Delta g_{\text{V}_{\text{O}}^{..}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}-\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$
eV
$\Delta g_{\text{dop}}$
The difference of formation energy of dopant between core and bulk,
i.e. $\Delta g_{\text{dop}}=\mu_{\text{dop},\text{c}}^{0}-\mu_{\text{dop},\text{b}}^{0}$
eV
$N_{\text{V}_{\text{O}}^{..},\text{b}}$ Number of available sites of oxygen vacancy in the bulk phase $\frac{1}{\text{m}^{3}}$
$N_{\text{V}_{\text{O}}^{..},\text{c}}$ Number of available sites of oxygen vacancy in the core region $\frac{1}{\text{m}^{3}}$
$N_{\text{dop},\text{b}}$
Number of available sites of dopant in the bulk phase.
$N_{\text{dop},\text{b}}=1/a^{3}$, with $a$ being the perovskite lattice constant
$\frac{1}{\text{m}^{3}}$
$N_{\text{dop},\text{c}}$ Number of available sites of dopant in the core region $\frac{1}{\text{m}^{3}}$
$z_{\text{V}_{\text{O}}^{..}}$ Valency number of oxygen vacancy, $z_{\text{V}_{\text{O}}^{..}}=2$ -
@@ 2.1 The Mott-Schottky model
The Mott-Schottky model assumes that the dopant is immobile and its concentration remains fixed and constant everywhere, i.e. $\nabla c_{\text{dop}}=0$. Note that this assumption simplifies the scenario but breaks the equilibrium condition of the electrochemical potential of the dopants. Based on this assumption, Poisson’s equation can be written as
@@
Substituting $c_{\text{dop},\text{b}}(\infty)=2c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$ into Eq. (12), we have
@@
With $l_{\text{D}}=\sqrt{\frac{\epsilon_{0}\epsilon_{\text{r}}k_{\text{B}}T}{2e^{2}c_{\text{dop},\text{b}}(\infty)}}$ being the Debye length and $\alpha=\frac{e}{k_{\text{B}}T}$, Eq.(13) becomes
@@ 2.2 The Gouy-Chapman model
The numerical solutions of the MS and GC models have been achieved through an iterative process de2009formation . In the present work, we propose an alternative numerical method via finite element method (FEM). By considering the symmetric configuration of a bicrysal (ignoring the grain misorientation), the semi-infinite simulation domain of $0,\infty)$ is sufficient. The location of the GB core is located at the left boundary of the semi-infinite domain ($x=0$). As for the boundary conditions, the electrostatic field is grounded at the right side, i.e., $\phi|_{x=\infty}=0$, while at the left side, a mixed boundary condition is given by
@@
where $Q_{\text{c}}(\phi_{0})=ew_{\text{c}}(z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{c}}+z_{\text{dop}}c_{\text{dop},\text{c}})$ and $\phi_{0}=\phi|_{x=0}$. We employed the boundary element in the finite element method to regard directly the mixed type of boundary condition defined in the above equation. The obtained numerical solutions agree fully with these from Ref. de2009formation (See Supplementary).
@@ 3.1 Free energy density functional and electrochemical potentials
In this part, we aim to develop a defect-chemistry informed phase-field model including not only the dominant driving force of SCL formation, i.e., segregation energy of oxygen vacancy, but also the segregation energy of acceptor dopant. The segregation energy of acceptor dopant can not govern the SCL formation, whereas it plays an important role on GB potential usler2024point , which is barely investigated by phase-field simulations reported in previous literature. In addition, we also include different available number sites of oxygen vacancy and acceptor dopant in bulk and GB core, i.e. $N_{\text{V}_{\text{O}}^{..},\text{b}}\neq N_{\text{dop},\text{b}}$ and $N_{\text{V}_{\text{O}}^{..},\text{c}}\neq N_{\text{dop},\text{c}}$.
Assuming an Fe-doped STO polycrystalline system consisting of $n$ grains, we choose free energy based Kim-Kim-Suzuki (KKS) phase-field model and utilize a set of non-conserved order parameters (OPs) $\eta_{i}$} to distinguish between different grains kim1999phase . Then, by treating the GB core and bulk as distinct phases with separate segregation energies cha2002phase , the phase-field model presented here aligns with the assumptions of the abrupt GB||SCL model (see Section 2) and inherently incorporates solute drag effects. Therefore, within grain $i$, we have $\eta_{i}=1$ and $\eta_{j}=0$ for $j<n$ and $j\neq i$. Within the GB core region of grain $i$ and $j$, the order parameters $\eta_{i}$ and $\eta_{j}$ spatially vary from 0 to 1. In the latter context, $\eta$ denotes the entire set of OPs $\eta_{i}$}. When considering electrochemical contributions, we employ two conserved concentration fields along with an electrostatic potential field. Other normalized field variables are $C_{\text{V}_{\text{O}}^{..}}$ for oxygen vacancy concentration, $C_{\text{dop}}$ for dopant concentration, and $\phi$ for the electrostatic potential. In Fe-doped STO, $\text{Fe}^{3+}$ and $\text{Fe}^{4+}$ can coexist. At high temperatures and low oxygen partial pressures, a higher proportion of $\text{Fe}^{3+}$ is observed and thus the dopant $\text{Fe}_{\text{Ti}}^{{}^{\prime}}$ has negative charge. For example, during the sintering process at 1623 K in air, approximately 90% of the iron is estimated to be in the $\text{Fe}^{3+}$ state, according to the bulk defect chemical model (see Fig. A1 in Ref. merkle2020water ). Thus, in this study, we assume the valence state of Fe to be 3+ and therefore the valance of dopant $\text{Fe}_{\text{Ti}}^{{}^{\prime}}$ is -1, i.e. $z_{\text{dop}}=-1$. Then, the formulation of total free energy density functional is given by
@@
The free energy density functional contains three contributions. The first term $f^{\text{grad}}$ represents the gradient energy of the diffuse interface and is given by
@@
where $\kappa$ is the coefficient of the gradient term. The second term $f^{\text{loc}}$ is the multi-well potential and it reaches minimum value in the bulk phase. According to Ref. moelans2008quantitative , aagesen2024electrochemical , its formulation is given by
@@
with $\omega$ being the free energy barrier coefficient, $\gamma=1.5$ for producing symmetric profile of $\eta$ and isotropic grain growth.
The determination of parameters such as $\kappa$ and $\omega$ can be linked to characteristics such as GB core width ($w_{\text{c}}$) and GB energy ($\sigma$) moelans2008quantitative . Their relations can be expressed as
@@
In the present work, we only consider the isotropic GB energy, i.e. it is not dependent on GB misorientation. The third term in free energy functional is the electrochemical contribution. Here, we treat the electrochemical free energy density as a mixture of bulk phase and GB core by an interpolation function. Thus, with the consideration of electrostatic free energy according to Refs. bishop2003effect , garcia2004thermodynamically , $f^{\text{ech}}$ is
@@
where the subscripts b and c indicate the bulk phase and GB core, respectively. $h(\eta)$ is the interpolation function and is given by
@@
Here, $h(\eta)$ exhibits non-monotonic behavior. Specifically, for an $i/j$ GB, $h(\eta_{i})=0$ when $\eta_{i}=1$ (representing the bulk phase), whereas $h(\eta_{i})=1$ when $\eta_{i}=0.5$ (representing the core). This interpolation function allows us to differentiate between the bulk phase and the core region. $\epsilon_{0}$ is the vacuum permittivity; $\epsilon_{\text{r}}$ is the relative permittivity. For STO, $\epsilon_{\text{r}}=9\times 10^{4}/(T-35)$ de2009formation . $\epsilon_{\text{r}}$ is only dependent on temperature in the present work. Hence, $\epsilon_{\text{r}}$ is constant across the SCL zone for specific temperature. Furthermore, $f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})$ and $f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})$ in Eq. 23 denote the electrochemical free energy contributions from the bulk and core, respectively, with $C_{\text{V}_{\text{O}}^{..},\text{b}}$ and $C_{\text{dop},\text{b}}$ being the phase concentrations of oxygen vacancy and dopant in the bulk phase, while $C_{\text{V}_{\text{O}}^{..},\text{c}}$ and $C_{\text{dop},\text{c}}$ being the phase concentrations of oxygen vacancy and dopant in the core. When $h(\eta)=1$, $C_{\text{dop},\text{c}}$ and $C_{\text{V}_{\text{O}}^{..},\text{c}}$ correspond to $c_{\text{dop},\text{c}}$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}$. When $h(\eta)=0$, $C_{\text{dop},\text{b}}$ and $C_{\text{V}_{\text{O}}^{..},\text{b}}$ are same as $c_{\text{dop},\text{b}}(x)$ and $c_{\text{V}_{\text{O}}^{..},\text{b}}(x)$. Local concentration is treated as the mixture of phase concentrations in bulk and core. The relations between local concentration and phase concentration are expressed as
@@
In principle of defect chemistry thermodynamic model as mentioned in Section 2, the electrochemical free energies $f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})$ and $f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})$ deliver the required equilibrium thermodynamic information in bulk and GB core and their expressions are given by
@@
$\begin{split}f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}C_{\text{V}_{\text{O}}^{..},\text{c}}+g^{0}_{\text{dop},\text{c}}C_{\text{dop},\text{c}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{c}}\ln\left(\frac{C_{\text{V}_{\text{O}}^{..},\text{c}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}}\right)+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}})\ln\left(\frac{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}}\right)\right.\\ &\left.\left.+C_{\text{dop},\text{c}}\ln\left(\frac{C_{\text{dop},\text{c}}}{\tilde{N}_{\text{dop},\text{c}}}\right)+(\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}})\ln\left(\frac{\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}}}{\tilde{N}_{\text{dop},\text{c}}}\right)\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{c}}+z_{\text{dop}}C_{\text{dop},\text{c}}\right)\phi,\end{split}$ (27)